OpenStudy (anonymous):

How do i factor the polynomial x^2+4x+3.

OpenStudy (anonymous):

(x+3)(x+1)

OpenStudy (math&ing001):

Try to divide the second term (4x) according to the last term (3) like this : x² + 4x + 3 = x² + x + 3x + 3 = x(x+1) + 3(x+1) = (x+3)(x+1)

OpenStudy (anonymous):

Thx guys really helped :)

OpenStudy (math&ing001):

Welcome :3